Calculate one of each of the following questions created by 3 different classmates

a. Mean and standard deviation given , looking for the percentage between two x values
b. Mean and standard deviation given , looking for the percentage above a certain x value
c. Mean and standard deviation given , looking for the x value at a certain percentile
1
Expert's answer
2020-06-08T19:13:05-0400

a) let the two values of x be x1 and x2

P(x1<x<x2)

="\Phi(\frac {x_2-\mu}{\sigma})-\Phi(\frac {x_1-\mu}{\sigma})"

b)P(X>x)

="1-\Phi(\frac {x-\mu}{\sigma})"

c) we find "z=\Phi^{-1}(percentile)"

"x=z\sigma+\mu"